OpenStudy (anonymous):

quick help anyone help please There are two 3-digit numbers n having the property that n is divisible by 11 and n/11 is equal to the sum of the squares of the digits of n. Find both values of n. You may submit them in either order.

OpenStudy (anonymous):

@phi @Preetha @e.mccormick @SithsAndGiggles @TylerD @badatmath_01

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

n = 11k since n > 99 and n < 1000, k is between 10 and 90. (110 < n < 990) k thus has 2 digits, so let k = a*10 +b n = 11*k = 11*(a*10 + b) =a*110 + 11*b = a*100 + 10*a*b + 1*b assuming temporarily that a*b*10 < 100 (so that a*b is the second digit of n): k = a^2 + (a*b)^2 + b^2 11k = n, so: 11*(a^2 + (a*b)^2 + b^2) = a*100 + 10*a*b + 1*b letting b = 0, solving quadratic for a gives a = 5 so: k = 50 -> n = 550 for one solution, the only such solution that a*b < 10. So then, knowing that a*b > 10, we can deduce that b is greater than 1. We also note that the 10s digit of n must be zero, as b is less than 10. As such, the sum of the digits must therefore be 11, which leads us to the set: (2,9);(3,8);(4,7);(5,6). Squaring and adding gives: 85,73,65,61. multiplying by 11: 935,803,715,671 of which 803 is the obvious answer, being the only one containing 0, 3, and 8. Therefor, we have k = 50 and k = 73. giving: n = 550 and n = 803